OpenStudy (anonymous):

stumped on another one! 3sqrt x-4=2 solve algebraically

OpenStudy (anonymous):

cube both side, resulting in x-4=2^3, which then gives you: x-4=8. Finally add 4 to both sides giving you: x=12

OpenStudy (anonymous):

umath has it if you mean \[\sqrt[3]{x-4}=2\]

OpenStudy (anonymous):

if you meant \[3\sqrt{x-4}=2\] however the answer would be different. i assume you meant the first one

OpenStudy (anonymous):

Umath has the right equation. I am trying to figure it out righrt now to see if it works

OpenStudy (anonymous):

I missed one step on this one!

OpenStudy (anonymous):

\[(\sqrt[3]{x-4?})^{3}=2^{3}\]

OpenStudy (anonymous):

\[x-4=8\]\[x=8+4\]\[x=12\]

OpenStudy (anonymous):

check\[\sqrt[3]{12-4}=\sqrt[3]{8}\]\[\sqrt[3]{2*2*2}=2\]